### v9

@@ 5 Mechanics of free growth
Some growth is capital widening, where structures and implements increase in number but do not change in design. Capital widening, however, is practical only so far before glut and diminishing returns set in. Further growth from that point must come from capital deepening, meaning improvements in the design of capital. Solow (1956) noted a kind of middle ground between capital widening and capital deepening in the disembodied growth mentioned earlier; ships carrying coals to Newcastle can raise prospective cash flows, and hence present value, by reversing the business plan. But Solow, who came to conclusions similar to ours from different evidence, puzzled as to how capital growth without net saving could be possible for capital deepening through “embodied” growth, where products of new design are made from plant of new design. [^4] The terms capital deepening, capital widening, embodied growth and disembodied growth are all Solow’s.
The solution, we suggest, is that embodied growth is disembodied growth on a finer scale. At each step toward realization of the new plant and products, raw materials and products and labor skills and plant capacity currently available on the market are adapted to new uses. The innovator pays for these inputs at a market price determined by their value in established productive uses, but applies them innovatively to realize higher prospective cash flows, and hence higher present values, to the innovator (Marshall (1890), Schumpeter and Opie (1934)). This difference in present value realized less price paid will here be called the “innovator’s reserve”, meaning reserve price for inputs of capital and labor.555i.e., capital and labor inputs are worth more to the innovator in that the innovator applies them in ways to realize greater returns. The present value of additional cash flow enabled by this advantage in return quantifies the innovator’s reserve and equivalently the non-random component of free growth. The innovator’s reserve quantifies the part of free growth explained by productivity gain as distinct from random market noise. As such, it is the quantity added to depreciation saving to enable embodied growth, so that net saving is never needed.

### v10

@@ 4 Inferences for the creative destruction idea
5 Solow’s puzzle
Some growth is capital widening, where structures and implements increase in number but do not change in design. Capital widening, however, is practical only so far before glut and diminishing returns set in. Further growth from that point must come from capital deepening, meaning improvements in the design of capital. Solow (1956) noted a kind of middle ground between capital widening and capital deepening in the disembodied growth mentioned earlier; ships carrying coals to Newcastle can raise prospective cash flows, and hence present value, by reversing the business plan. But Solow, who came to conclusions similar to ours from different evidence, puzzled as to how capital growth without net saving could be possible for capital deepening through “embodied” growth, where products of new design are made from plant of new design.
Solow’s Nobel Prize acceptance speech in mentions the embodiment problem: [^4] Solow (1988). The terms capital deepening, capital widening, embodied growth and disembodied growth are all Solow’s.
"… much technological progress, maybe most of it, could find its way into actual production only with the use of new and different capital equipment. Therefore the effectiveness of innovation in increasing output would be paced by the rate of gross investment. A policy to increase investment would therefore lead to … a faster transfer of new technology into actual production … By 1958 I was able to produce a model that allowed for the embodiment effect. If common sense was right, the embodiment model should have fit the facts better than the earlier one. But it did not… I do not know if that finding should be described as a paradox, but it was at least a puzzle."
Solow was right in that much innovation requires months or years of gross investment to pay for negative cash flows in building new plant before the new products roll out from the factory doors. What he left unsaid is that this investment would not have been forthcoming unless positive cash flows at the end of the tunnel were predicted to be sufficient to create positive present value from the start. Data in our displays suggest that this initial creation of value, plus depreciation investment, are enough.
@@ 6 Mechanics of free growth
The present value argument just given offers an accounting explanation of embodied growth in terms of sources of funds. The step-by-step mechanical explanation, we suggest, in terms of things rather than dollars, is that embodied growth is disembodied growth on a finer scale. At each step toward realization of the new plant and products, raw materials and products and labor skills and plant capacity currently available on the market are adapted to new uses. The innovator pays for these inputs at a market price determined by their value in established productive uses, but applies them innovatively to realize higher prospective cash flows, and hence higher present values, to the innovator (Marshall (1890), Schumpeter and Opie (1934)). This difference in present value realized less price paid will here be called the “innovator’s reserve”, meaning reserve price for inputs of capital and labor.555i.e., capital and labor inputs are worth more to the innovator in that the innovator applies them in ways to realize greater returns. The present value of additional cash flow enabled by this advantage in return quantifies the innovator’s reserve and equivalently the non-random component of free growth. The innovator’s reserve quantifies the part of free growth explained by productivity gain as distinct from random market noise. As such, it is the quantity added to depreciation saving to enable embodied growth, so that net saving is never needed.
@@ 16 Discussion and conclusions
Solow’s puzzle runs deep. How can thoughts be things? How can assessments of present value pay, at least in part, for their own realization? Science and philosophy push our minds into areas where they are not at home. Neither relativity nor quantum mechanics, which explore the scale of great speeds and minute distances, sink in easily. Economics, too, invites surprises as it looks farther from the scale of the household and community and firm to which we are accustomed. Free growth theory predicts only at the universal scale, and accepts that thrift explains a greater part of growth at progressively smaller scales. Our findings, with that allowance, are awkward for much of economics as generally taught. We encourage others to check our data and methods and inferences, and improve on them wherever they can.